# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v1

@@ V-A EKF initialization
We initialize the state vector $\mathbf{x}$ to 0, attitude $\mathbf{q}$ to the unit quaternion, and covariance matrix $\mathbf{P}$ to 0 except for the following sub-blocks:
$\mathbf{P}_{b_{a}\,b_{a}}=\mathbf{I}\eta^{2}_{b_{a}};\;\mathbf{P}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\eta^{2}_{b_{\omega}};\;\mathbf{P}_{\delta\theta_{xy}\,\delta\theta_{xy}}=\mathbf{I}\eta^{2}_{\theta},$ (8)
where $\eta_{b_{a}}$, $\eta_{b_{\omega}}$ and $\eta_{\theta}$ are the initial uncertainties (standard deviations) of the accelerometer bias, gyroscope bias and roll/pitch attitude error. We consider the initial yaw attitude uncertainty to be zero because we are not incorporating a compass/magnetometer.
@@ V-B EKF propagation
$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (13)
$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (14)
@@
$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta\,\omega}=\mathbf{C}^{w}_{b}t;$ (17)
$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (18)
@@ V-C EKF update and error state reset
As previously mentioned, the attitude part of the robot’s pose is expressed as error state in the EKF, and after an EKF update the nominal value of $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
@@
where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for the given subblock.
@@ V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. The returned point cloud is first transformed by $\mathbf{C}_{r}^{b}$ to match the IMU orientation (we consider the separation between IMU and radar to be negligible). Afterwards, we use the RANSAC-LSQ algorithm [3] to both filter out outlier (dynamic) points, and estimate the current body-frame velocity of the robot, including a covariance matrix. The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
$\mathbf{r}=\mathbf{y}-\mathbf{C}^{b}_{w}\mathbf{v};\;\mathbf{H}_{:\,v}=\mathbf{C}^{b}_{w};\;\mathbf{H}_{:\,\delta\theta}=\mathbf{C}^{b}_{w}[\mathbf{v}]_{\times}.$ (23)
@@ V-E Scan matching observation
Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector $\mathbf{H}_{c}\mathbf{r}$, observation matrix $\mathbf{H}_{c}\mathbf{H}$ and covariance matrix $\mathbf{H}_{c}\mathbf{R}\mathbf{H}_{c}^{T}$ by a constraint matrix defined as follows:
$\mathbf{H}_{c}=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\end{bmatrix}.$ (24)
Table I: Quantitative Analysis on NTU4DRadLM Dataset using EVO. Best results in bold, second best results underlined
@@
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
NTU4DRadLM baseline: gicp [25] 4.87 0.0648 5.65 0.0257 6.73 0.0094 6.03 0.0097
NTU4DRadLM baseline: Fast-LIO [25] 3.97 0.0773 4.47 0.0273 8.38 0.0096 6.95 0.0105
Our method (GICP scan matching) 1.69 0.0323 4.50 0.0198 3.92 0.0044 3.81 0.0054
Our method (single hypothesis) 1.47 0.0284 3.17 0.0150 2.95 0.0043 4.80 0.0065
Our method (full version) 1.33 0.0260 2.92 0.0134 3.49 0.0047 4.55 0.0061
Table II: Quantitative Analysis on NTU4DRadLM Dataset using rpg_trajectory_evaluation. Methods marked with * are self-reported metrics
cp nyl loop2 loop3
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
@@
NTU4DRadLM baseline: gicp [25] 4.921 0.068 5.775 0.027 6.743 0.010 6.242 0.010
NTU4DRadLM baseline: Fast-LIO [25] 4.178 0.083 4.759 0.030 8.221 0.010 7.120 0.011
Our method (GICP scan matching) 1.746 0.035 4.142 0.020 3.902 0.005 3.781 0.006
Our method (single hypothesis) 1.501 0.030 3.133 0.015 2.851 0.004 4.529 0.007
Our method (full version) 1.383 0.028 2.934 0.014 3.413 0.005 4.259 0.006
Figure 3: Generated trajectories for the 4 tested NTU4DRadLM sequences.
@@ VI Experimental results
We compare our odometry with previous published 4D radar-inertial odometry methods on the NTU4DRadLM dataset [25]. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [26], as well as other works such as EFEAR-4D [22] or RIV-SLAM [20]. We compare versions of these methods without loop closure if possible, in order to focus on evaluating pure odometry. Out of the six sequences present in NTU4DRadLM, we perform evaluation using four: cp and nyl (low speed regime, using a handcart platform), and loop2 and loop3 (high speed regime, using a car platform). The garden and loop1 sequences cannot be reliably used for evaluation because they contain large gaps in the IMU and radar data, respectively.
We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [6], as well as the older rpg_trajectory_evaluation package [27] used by previous works, using default settings. In the case of evo we use a custom evaluation script written specifically to follow the default behavior of rpg_trajectory_evaluation as closely as possible. We run both evaluation packages on the baseline GICP and Fast-LIO trajectories published along with the NTU4DRadLM dataset, as well as the trajectories generated by our odometry solution. In particular, we report results for the full version of our method, and also two ablated versions: one that replaces the Gaussian modeling and scan matching with GICP, and another that uses single-hypothesis scan matching. The GICP version uses small_gicp [9] as the underlying implementation, and in both ablations we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm). We also include the self-reported metrics for all methods that appear in their respective papers as-is, since the trajectory files needed for metric calculation are not available.
@@
The source code of our odometry solution can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio. We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo. The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, number of Gaussians, etc.
VI-A Quantitative Analysis
VI-A1 Evaluation with evo
In Table I, we can observe major improvements in all 4 sequences provided by our method compared to the baselines originally tested by the authors of NTU4DRadLM. Moreover, the comparison between the Gaussian versions of our method and our GICP ablation shows improvements in relative translation error ($t_{rel}$) in 3 out of 4 sequences, and improvements in relative rotation error ($r_{rel}$) in 3 out of 4 sequences. In particular, the best improvements can be seen in the nyl sequence. The version with scan matching based on multiple hypotheses overperforms the version with a single hypothesis in 3 out of 4 sequences. We theorize that the sequences where our GICP ablation overperforms Gaussian versions of our method contain more challenging/sparser geometry due to being higher speed car driving sequences in open environments. Nonetheless, we believe it should be possible to further adjust the hyperparameters used (such as number of Gaussians used for modeling, number of hypotheses used for scan matching, keyframe thresholds, EKF covariances used, etc.) in order to optimize the algorithm for those sequences.
VI-A2 Evaluation with rpg_trajectory_evaluation
In Table II, we note discrepancies in the NTU4DRadLM baseline metrics between the reported values in the paper and the values obtained by running the evaluation software on the corresponding published trajectories, especially in the higher speed sequences. We also note very similar results for all methods generated by both evaluation packages, validating the use of both packages for evaluating 4D radar-inertial odometry methods. The errors reported by EFEAR-4D for the higher speed sequences are much higher than the values reported by other methods. We believe this to be caused by some difference in the evaluation methodology, as the same paper reports similar errors for baseline methods such as [26]. The most challenging sequence is loop3, and interestingly the best results are the ones reported by the NTU4DRadLM authors for the simplest baseline method. RIV-SLAM is the best performing method in loop2, however it must be said that to the best of our knowledge this method includes loop closure. The second best performing method in that sequence is the single-hypothesis version of our method, in line with the results in Table I.
VI-B Qualitative Analysis
Figure 3 shows the plotted trajectories of the ground truth, NTU4DRadLM GICP baseline and the full version of our method; across the 4 sequences. We can observe that in the low speed sequences (cp, nyl) our method accumulates substantially less drift in the XY plane, and manages to closely match the ground truth. High speed sequences (loop2, loop3) prove to be more challenging for both methods; in particular, our method drifts the most in loop3 with respect to the ground truth. On the other hand, our method performs much better than the baseline in the XZ plane across all 4 sequences. In particular, in loop2 and loop3 our method avoids large drifts in the Z axis, compared to the drifts in the baseline of over 70 meters. We think the higher accuracy in the XZ plane for our methods is responsible for improved quantitative results in these sequences, despite the lower accuracy in the XY plane.
@@ VI-C Summary of results
We consider that the results described above validate our methodology. Our methods produce better relative translation and rotation metrics in 3 out of 4 tested sequences. Moreover, they produce generally better results regarding Z axis drift. Our ablation experiments also validate the Gaussian-based modeling and scan matching algorithm, as well as the multi-hypothesis approach; which contribute to improve both estimation accuracy and robustness.

### v2

@@ V-A EKF initialization
First, we initialize all components of the state vector $\mathbf{x}$ to 0, and the rotational variables to the unit quaternion. The initial radar-to-IMU frame ($\mathbf{t}_{r}^{b}$, $\mathbf{q}_{r}^{b}$) can be supplied externally. The robot is assumed to remain still during the first few seconds, which allows for averaging initial accelerometer/gyroscope readings and extracting an estimate of the body-frame gravity vector $\tilde{\mathbf{g}}_{b}$ (after flipping sign) and the gyroscope bias $\mathbf{b}_{\omega}$ respectively. An initial value for the roll/pitch components of $\mathbf{q}_{b}^{w}$ can then be derived from $\tilde{\mathbf{g}}_{b}$, while the yaw component is set to 0 since we are not incorporating a compass/magnetometer. In addition, we derive the initial accelerometer bias as $\mathbf{b}_{a}=\mathbf{C}_{w}^{b}\left(\mathbf{C}_{b}^{w}\tilde{\mathbf{g}}_{b}-\mathbf{g}\right)$, thus taking into account the difference in magnitude between the expected and observed gravity vectors.
We initialize the covariance matrix mostly as 0, except for the subblocks related to $\mathbf{t_{r}^{b}}$, $\mathbf{b}_{a}$, $\mathbf{b}_{\omega}$, $\delta\boldsymbol{\theta}_{b}^{w}$ and $\delta\boldsymbol{\theta}_{r}^{b}$, whose diagonals we fill in according to supplied initial values.
@@ V-B EKF propagation
$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta_{b}^{w}}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (8)
$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta_{b}^{w}\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta_{b}^{w}}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (9)
@@
$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta_{b}^{w}\,\omega}=\mathbf{C}^{w}_{b}t;$ (12)
$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta_{b}^{w}\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (13)
@@ V-C EKF update and error state reset
As previously mentioned, rotations are expressed as error state in the EKF, and after an EKF update the nominal value of each $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
@@
where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for every $\delta\theta\,\delta\theta$ subblock.
Finally, in order to prevent especially noisy observations from causing instability in the filter, we perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution, like [4].
@@ V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. We use the RANSAC-LSQ algorithm [4] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix. The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
$\mathbf{r}=\mathbf{y}-{\mathbf{C}_{b}^{r}\left([\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}\right)};$ (18)
$\mathbf{H}_{:\,v}={\mathbf{C}_{b}^{r}}\mathbf{C}_{w}^{b};\;{\mathbf{H}_{:\,t_{r}^{b}}=\mathbf{C}_{b}^{r}[\boldsymbol{\omega}]_{\times}};\;{\mathbf{H}_{:\,b_{\omega}}=\mathbf{C}_{b}^{r}[\mathbf{t}_{r}^{b}]_{\times}};$ (19)
$\mathbf{H}_{:\,\delta\theta_{b}^{w}}={\mathbf{C}_{b}^{r}}\mathbf{C}_{w}^{b}[\mathbf{v}]_{\times};\;{\mathbf{H}_{:\,\delta\theta_{r}^{b}}=\mathbf{C}_{b}^{r}[[\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}]_{\times}}.$ (20)
@@ V-E Scan matching observation
Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined.
Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching
@@
Usage $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
IMU EV SM (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
@@
EKF-RIO [4] 14.76 0.3955 35.89 0.2464 44.31 0.0349 50.23 0.0910
(Ablation) Our EKF with NDT [2] 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
(Ablation) Our EKF with GICP [21] 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
(Ablation) Our EKF with VGICP [12] 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
(Ablation) Gaussian single hypothesis 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
Gaussian multi-hypothesis 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
Table II: Results on Snail-Radar Dataset (Oculii Eagle). Best results in bold, second best results underlined.
st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
(Ablation) Our EKF with NDT [2] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
(Ablation) Our EKF with GICP [21] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
(Ablation) Our EKF with VGICP [12] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
(Ablation) Gaussian single hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
Gaussian multi-hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
Table III: Results on Snail-Radar Dataset (ARS548). Best results in bold, second best results underlined.
st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
(Ablation) Our EKF with GICP [21] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
(Ablation) Our EKF with VGICP [12] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
(Ablation) Gaussian single hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
Gaussian multi-hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
@@ VI Experimental results
In order to evaluate the performance of our Gaussian modeling and scan matching method, we use existing public 4D radar datasets. In particular, we selected the NTU4DRadLM [29] and Snail-Radar [8] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [7], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible. In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m because the errors are much smaller.
Besides our full method, we evaluate four ablations within our EKF framework: one without multi-hypothesis registration, and three replacing our Gaussian-based algorithms with NDT, GICP and VGICP. Since these are all single-hypothesis methods, we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm). The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [11] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching.
@@
The source code of our odometry solution can be accessed online on GitHub222 https://github.com/robotics-upo/gaussian-rio-cpp . We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo. The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, etc.
VI-A Results on NTU4DRadLM
We compare our odometry with previous published results on the NTU4DRadLM dataset [29], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [30], as well as other works such as EFEAR-4D [26] or RIV-SLAM [24]. We include these self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public. Even though there are no published results for EKF-RIO [4] on this dataset, it is one of the few methods with available source code, and so we are able to produce and include evaluation results for it. We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
We evaluate the cp, nyl (low speed handcart), loop2 and loop3 (high speed car) sequences. The garden and loop1 sequences are excluded because they contain large gaps in the IMU and radar data, respectively.
Table I shows the comparison between all methods. We can observe how our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation). Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
VI-B Results on Snail-Radar
In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [8], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions. In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform. These sequences are all fairly long, with the longest (iaf_20231213_2) being 15 min and most being around 7 min in length.
Table II shows results using the Oculii Eagle radar. In general, we can observe the best results to be spread across the different methods. Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT). The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences. It is worth mentioning that NDT, unlike in NTU4DRadLM, seems to behave much better in Snail-Radar.
Table III shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be caused by the radar lacking the scan enhancement algorithm used by Oculii Eagle. However interestingly, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause the program to abnormally terminate, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
@@ VI-C Summary of results
We consider that the results described above validate our methodology. Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments. Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results. Our EKF-based radar-inertial framework is also validated by the experiments, which show how on a publicly evaluated dataset we are able to improve on the state of the art even when not using our Gaussian algorithms. On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
As an additional observation, it is worth mentioning that we believe the radar-to-IMU calibration file provided by NTU4DRadLM to be inaccurate, as it does not seem to match the placement of the different sensors in the photo they provide of their platform. The effects of this inaccurate calibration can be especially seen in the rotational errors, which are much higher for all methods compared to those achieved under Snail-Radar sequences.